Grade 11General Physics

a particle moves in positive x-direction according to law s=12t-t^2. Where t is time in second. Find average speed from t=0 to t=8sec is

Jagjeet Kaur
8 Years agoGrade 11

1 Answer

Venkat
8 Years ago
First calculate the distance travelled at 0 and 8 secs. Then calculate the average speed.
s = 12t-t2
s(0) = 0 m
s(8) = 12(8)-82 = 32 m
Average speed = (32-0)/8 = 4 m/s